### v1

# Undesirable Memorization in Large Language Models: A Survey

Abstract
While recent research increasingly showcases the remarkable capabilities of Large Language Models (LLMs), it’s vital to confront their hidden pitfalls. Among these challenges, the issue of memorization stands out, posing significant ethical and legal risks. In this paper, we presents a Systematization of Knowledge (SoK) on the topic of memorization in LLMs. Memorization is the effect that a model tends to store and reproduce phrases or passages from the training data and has been shown to be the fundamental issue to various privacy and security attacks against LLMs.
We begin by providing an overview of the literature on the memorization, exploring it across five key dimensions: intentionality, degree, retrievability, abstraction, and transparency. Next, we discuss the metrics and methods used to measure memorization, followed by an analysis of the factors that contribute to memorization phenomenon. We then examine how memorization manifests itself in specific model architectures and explore strategies for mitigating these effects. We conclude our overview by identifying potential research topics for the near future: to develop methods for balancing performance and privacy in LLMs, and the analysis of memorization in specific contexts, including conversational agents, retrieval-augmented generation, multilingual language models, and diffusion language models.
Index Terms:
Memorization, Large Language Models, Privacy in LLMs
I Introduction
In recent years, large language models (LLMs) have demonstrated remarkable advancements, driven by the scaling of model parameters, large amounts of data, or training paradigms [1, 2, 3, 4]. State-of-the-art models have exhibited capabilities across a broad spectrum of Natural Language Processing (NLP) tasks, consistently pushing the benchmark records in areas such as text generation, code synthesis, machine translation, question answering, and summarization [5, 6]. These models use self-supervised learning on massive datasets, enabling them to perform competitively or even surpass human-level performance in specific tasks [7, 8].
Despite these impressive advancements of LLMs, researchers have pointed out the problematic features of these models, including hallucination [9], bias [10], and privacy and security [11]. In the context of data privacy, memorization is one of the most significant issues. Memorization in LLM refers to the model’s tendency to store and reproduce exact phrases or passages from the training data rather than generating novel or generalized outputs. While memorization can be advantageous in knowledge-intensive benchmarks, such as factual recall tasks or domain-specific question answering [12], it introduces significant ethical and legal challenges: First, models may inadvertently reveal sensitive or private information included in their training data, posing significant privacy and security risks [13, 14]. Second, the ability of LLMs to repeat verbatim copyrighted or proprietary text from their training data raises issues related to intellectual property infringement [15, 16].
These challenges motivate the need to further explore memorization in LLMs to effectively tackle its associated challenges. We provide an overview of aspects related to memorization, emphasizing the need for further exploration of this topic, thereby balancing model performance with the risks of privacy breaches and ethical concerns.
Before recent advances in LLMs, memorization was explored extensively as a topic in machine learning and deep learning. Usynin et al. [17] explore memorization in machine learning. They propose a framework to quantify the influence of individual data samples and detect memorization in various learning settings. Wei et al. [18] provide a systematic framework for understanding memorization in deep neural networks, discussing LLM memorization as a type of deep neural network. Survey papers on the privacy and safety of LLMs often address memorization as a core phenomenon, framing it as both a privacy issue and a foundational factor that supports other challenges they explore [19, 20]. Hartmann et al. [21] provide an overview of memorization in general-purpose LLMs. They aggregate memorization-related topics from the copyright, privacy, security, and model performance perspectives. In our survey, we not only add more recent work, but we specifically focus on memorization as an undesirable phenomenon, with the aspects intentionality, severity of memorization, retrievability, and transparency of memorization. In addition, we provide an extensive and concrete research agenda for the near future. Table I provides a comparison of our scope to the scopes of previous surveys.
Paper Focus Differences With our Study
Usynin et al. [17] Memorization in general machine learning We focus on LLMs
Wei et al. [18] Memorization in deep neural networks We focus on LLMs
Neel and Chang [19] Privacy Problems in LLMs Memorization is treated as an instance of a privacy related problems rather than being the focus.
Smith et al. [20] Privacy Problems in LLMs Memorization is treated as an instance of a privacy related problems rather than being the focus.
Hartmann et al. [21] Memorization in general-purpose LLMs We focus on unintended and undesirable memorization in LLMs, and propose concrete directions for future work.
TABLE I: List of existing surveys vs our work
In the following sections of this paper, we provide a comprehensive exploration of memorization in LLMs, structured into six main sections. We begin with the Spectrum of Memorization[II], where we examine the concept from multiple perspectives including intentionality, retrievability, severity, abstraction, and transparency, offering a deep understanding of how memorization occurs. Next, in Measuring Memorization [III], we review various methodologies that have been used in previous studies to quantify and assess memorization within LLMs. We then explore the Influences and Dynamics of Memorization [III], identifying key factors and conditions that contribute to this phenomenon. Following this, in section [V] we analyze how memorization appears in specific LLM architectures. Next, in the Mitigating Memorization section [VI], we discuss strategies and techniques employed to minimize or control memorization in LLMs, addressing concerns related to privacy and generalization. Finally, we conclude with Future Directions and Open Challenges [VII], outlining potential areas for further research and unresolved questions, before summarizing our findings in the Conclusion [VIII]. Figure 1 shows a visual overview of our systemization scope.
Figure 1: Graphic summary of our systematization of knowledge
II Spectrum of Memorization
II-A Intentionality
Although memorization of factual information can be helpful for the model the model perform more accurately on benchmark tasks such as question answering, other data might be retained without any clear purpose, and this might cause issues with privacy and copyright. This raises questions about whether certain instances of memorization is deliberate (by design) or incidental and undesirable. Understanding the difference between deliberate memorization and unintended retention is crucial for both improving model design and addressing issues like privacy and data security [22, 23, 24].
Unintended memorization
Carlini et al. [25] explores the concept of unintended memorization in neural networks, identifying it as a form of memorization that occurs independently of overfitting and often early in the training process, before the learning has fully converged. To understand and measure this phenomenon, they introduced the “exposure metric” as part of a testing methodology (see Section III-A). It is important to note that the language modeling loss function itself is chosen in a way to be optimal when the language model’s output are recreations of training data, however, for the unintended memorization, this happens even when the loss function still has some distance to its optimal value. Unintended memorization could lead to various issues including:
  • Privacy risks: LLMs might inadvertently memorize and potentially reveal sensitive personal information present in the training data [16].
  • Security vulnerabilities: Unintended memorization of confidential information like API keys or passwords could lead to security breaches [25].
  • Copyright violation: Memorization of copyrighted material may lead to legal challenges, especially if the model can reproduce substantial portions of protected works [16].
Deliberate memorization
Deliberate memorization, in contrast, refers to the intentional design of LLMs to retain specific types of information from their training data. This is often a desirable feature, as it allows models to store and utilize vast amounts of knowledge. Key aspects of deliberate memorization include:
  • Knowledge retention: LLMs are often deliberately designed to memorize facts, concepts, and general knowledge to enhance their performance in tasks like question answering and information retrieval [26].
  • Language generation: Deliberate memorization of linguistic patterns, grammar rules, and vocabulary is crucial for the model’s ability to generate coherent and contextually appropriate text.
  • Alignment goals: Deliberate memorization can be utilized in the AI alignment phase to inject desired behaviors and values into the model [27].
As Ranaldi et al. [28] show in their experiments, deliberate memorization is generally beneficial for model performance, however, it also raises important considerations:
  • Bias and fairness: The choice of what information to deliberately memorize can introduce or amplify biases in the model output.
  • Transparency and auditing: Understanding what has been deliberately memorized is crucial for model interpretability and auditing purposes.
II-B Severity of memorization
While memorization in LLMs is widely discussed, there is no universally accepted definition. The definition of memorization can vary based on factors such as the level of detail required, the context in which the information is recalled, and the specific task or application at hand. This section discusses different severity levels of recall that are explored in the literature evolving around memorization in LLMs, including perfect memorization, verbatim memorization, and approximate memorization.
Perfect memorization
Perfect memorization refers to a scenario where a model retains and can reproduce all training data exactly. Kandpal et al. [29] define Perfect Memorization as follows:
Definition (Perfect Memorization).
A model is said to exhibit perfect memorization behavior if it only assigns non-zero probability to the samples it has been trained on. Sampling from a perfect memorization model is identical to sampling from the training data.
This form of memorization acts as an imaginary language model and is used as a baseline for the experiments.
Verbatim memorization
Verbatim memorization is a more specific form of memorization that involves the exact reproduction of strings from the training data. It captures instances where the model outputs a training sample without any alterations. This type of memorization is straightforward to identify and is often the primary focus in discussions about memorization in language models [30, 13].
Carlini et al. [13] define verbatim memorization under the “Eidetic Memorization” terminology as follows:
Definition (Verbatim memorization).
a string $s$ from the training set is defined as verbatim memorized if there exists a prompt $p$ so that $LM(p)=s$.
Tirumala et al. [31] consider another version of this as “Exact memorization”:
Definition.
[Exact memorization] a context $c=(p,s)$ from the training set is being memorized if $Argmax(LM(p))=s$.
This definition is more strict and does not capture the memorization under different forms of sampling methods, rather it only considers the greedy sampling [IV-E].
It should be noted that these definitions do not impose any restrictions on the length of prompt $p$ or the generation $s$. However, if the generated text is too short, it may not be appropriate to classify it as memorization. Typically, a certain number of tokens is considered when assessing memorization; for example, [23, 31] use a 50-token output restriction.
Approximate memorization
Approximate memorization extends beyond verbatim memorization to include instances where the output is similar but not identical to the training data. Verbatim memorization does not capture the subtler forms of memorization, as it is too confined. For example, if two sentences differ by a minor detail, such as punctuation, a misspelled word, or a stylistic variation (e.g., American English vs. British English), these instances would fall outside the strict boundaries of verbatim memorization. However, human judges would likely consider these variations as memorized examples.
Ippolito et al. [32] define “Approximate memorization” as follows:
Definition (Approximate memorization).
a suffix $s$ for prefix $p$ is labeled as approximately memorized if for generation $g=LM(p)$, $BLEU(g,s)>0.75$.
They discuss that this threshold was chosen based on qualitative analyses of the samples. Examples of this are shown in Table 2. By adopting this definition of memorization, they show that the measurement of memorization can increase by a factor of two. However, they also note that this definition can lead to both false positives and false negatives when compared to human judgment, indicating a potential direction for future investigations.
Figure 2: BLEU similarity score of more than 0.75 (from Ippolito et al. [32])
II-C Retrievability
The common way to analyze memorization in LLMs is through sampling. This involves selecting tokens probabilistically from the model’s predicted distribution at each step, to see if a sequence of probabilistically generated tokens can be traced back to the training data. Since the LLM’s input and output domains are both discrete, it is important to note that it might not be possible to extract all of the memorized content through given a finite number generation trials through sampling. Based on how we can make an LLM output the training data we can classify memorization into extractable and discoverable.
Extractable memorization
Extractable memorization refers to the ability to retrieve specific information from a model’s training data without direct access to that data. Carlini et al. [23] defines “Extractable memorization” as follows:
Definition.
(Extractable Memorization) Given a model with a generation routine $Gen$, an example $s$ from the training set $S$ is extractably memorized if an adversary (without access to $S$) can construct a prompt $p$ that makes the model produce $s$ (that is, $Gen(p)=s$) .
A more specific form of this is k-extractable memorization:
Definition.
(K-extractable Memorization) A string $s$ is said to be k-extractable if it (a) exists in the training data, and (b) is generated by the language model by prompting with $k$ prior tokens [33].
Analyzing extractable memorization usually involves two main challenges: Designing prompts that best elicit memorization in a model and verifying if the model output is indeed from the training data.
Research in this area has employed various strategies. Carlini et al. [13] recover training examples from GPT-2 by prompting it with short strings from the public Internet and manually verifying the outputs via Google search. This method confirmed the memorization of about 0.00001% of GPT-2’s training data due to the labor-intensive verification process. Nasr et al. [34] conduct extensive analysis on Pythia, RedPajama, and GPT-Neo models [35]. They query these models with millions of 5-token blocks from Wikipedia and count for unique 50-grams that the model generates whether they exist in a combined dataset of the models’ training data. Their method was more successful, showing that 0.1% to 1% of the models’ outputs are memorized, with a strong correlation between model size and memorization abilities.
Discoverable memorization
Discoverable memorization measures the extent to which models can reproduce their training data when explicitly prompted with data from their training set. Nasr et al. [34] suggests the following definition for “Discoverable memorization”:
Definition.
[…]
Discoverable memorization provides an upper bound for data extraction, compared to extractable memorization’s lower bound. Ideally, discoverable memorization requires querying the model with its entire training set, which is computationally intractable. Also, noteworthy is that discoverable memorization differs from extractable memorization in that the prompt $p$ is known to be from the training set.
Carlini et al. [23] investigate the upper bounds of data extraction in GPT-Neo models through discoverable memorization. They find that (a) LLMs discoverably memorize roughly 1% of their training datasets; (b) There is a log-linear correlation between data extraction and model size, repetition of data, and prefix context length. Other studies on different models (PaLM, MADLAD-400) corroborate the 1% memorization rate when prompting with about 50 tokens of context [36, 37].
Comparing discoverable and extractable Memorization
Nasr et al. [34] compare their extractable memorization results with the discoverable memorizations from Carlini et al. [23] for the GPT-Neo 6B parameter model. This comparison revealed that (1) some sequences are both discoverably and extractably memorized; (2) some sequences are discoverably memorized but not extractably memorized, and vice versa; (3) the overlap between these two types of memorization provides insights into the model’s information retention and retrieval mechanisms. This comparison highlights the complementary nature of these two approaches in understanding a model’s memorization capabilities and the retrievability of information from its training data.
II-D Abstraction
Abstraction in the context of memorization refers to the level of specificity at which information is retained and reproduced by language models. This spectrum ranges from highly specific, fact-based memorization to more generalized concept-level retention. This is related to the concept of verbatim memorization in Section II-B, but abstraction focuses more on information (semantics of facts) than on literal word use. In terms of abstraction, memorization can be categorized into two main types: factual memorization and conceptual memorization. These categories represent different levels of information granularity and have different implications for model performance, privacy concerns, and potential applications.
Factual memorization
Factual memorization in language models involves accurately retaining and reproducing specific pieces of information from training data, often relying on context cues to recall details such as names, dates, and numbers or the relations between such entities. Petroni et al. [12], Jiang et al. [38], AlKhamissi et al. [39], Luo et al. [40] investigate the measurement of factual knowledge in LLMs by evaluating their comprehension of facts from knowledge bases using ranking metrics. Pezeshkpour [41] proposes two approaches of explicit and implicit instillation of factual knowledge in LLMs while proposing to use KL-divergence as a metric for measuring factual memory. Yu et al. [42] show that the frequency of facts in the pretraining data significantly influences a language model’s behavior. They find that models are more likely to generate memorized capital cities for frequently mentioned countries, while being less responsive to in-context information. Additionally, they demonstrate that larger models tend to favor memorized answers over in-context details, even when the memorized facts are less frequent. Kang and Choi [43] also show that factual knowledge probing accuracy in large language models is linked to subject-object co-occurrence. They reveal that LLMs often prioritize frequently co-occurring word pairs, which can override correct answers, particularly when those answers are infrequent. Their findings highlight the need to address co-occurrence bias in future model development.
While factual memorization could be beneficial for the performance of LLMs in tasks namely question answering, it could potentially lead to various issues such as undesired bias and privacy and copyright violations.
Conceptual memorization
Conceptual memorization involves the model’s ability to internalize and generalize broader ideas, patterns, and concepts from its training data [44, 45, 46, 4]. That is, Lee et al. [47] explore the idea plagiarism of GPT-2 model and empirically show that language models go beyond merely generating verbatim text from their training data; rather they also rephrase sentences and imitate ideas from different sources.
Compared to factual memorization, conceptual memorization generally poses a lower risk of exposing specific, sensitive information. At the same time, it is more difficult to be detected and it may lead to perpetuation of biases present in the training data at a more systemic level.
The boundary between conceptual memorization and the understanding of the model (generalization) is subtle and still debated.
III Measuring memorization
Measuring memorization in large language models is a complex task, primarily because we cannot easily compute the probability of all the samples present in the training data. To obtain an accurate measure of memorization in language models, we would ideally need to compare the model’s output probabilities for every possible sequence against the training data. However, given the vast space of possible sequences, this is computationally intractable. Therefore, researchers have developed various methods to approximate and quantify memorization.
III-A Exposure metric
The Exposure Metric, introduced by Carlini et al. [30], provides a quantitative measure of how much a model has memorized specific sequences from its training data. This metric is particularly useful for assessing the memorization of rare or unique information.
The exposure metric has been widely adopted in memorization studies. To practically apply the exposure metric, researchers often use the “canary extraction test” [30]. This involves inserting known ‘canary’ sequences into the training data and then measuring their exposure in the trained model.
Helali et al. [48] propose the so-called “d-exposure” metric as a measure of memorization for discriminative tasks such as text classification, sinse in those situations we don’t have access to the perplexity of a given text.
III-B Inference attacks
Inference attacks are another approach to measuring memorization, focusing on the model’s ability to reveal information about its training data. These attacks typically perform under the assumption that when a model displays high confidence in its outputs, those outputs are likely to be from the training data, which means the model has memorized them. These attacks can be categorized into two main types:
Membership Inference Attacks (MIA)
These attacks aim to determine whether a specific data point was part of the model’s training set. Shokri et al. [49] introduced this concept for machine learning models, and it has since been adapted for language models. While the MIA approach in the context of machine learning uses a “shadow learning” paradigm, this is not feasible for LLMs, since they are much heavier in size. Instead, methods for MIA on LLMs try to extract the certainty of the LLMs on the given text and use it as sign of membership in the training data, which is usually done by computing the perplexity metric on the sequence [50, 51, 52, 53, 54, 55, 56].
Extraction Attacks
These attacks attempt to extract specific pieces of information from the model that were present in its training data. Different works have demonstrated the feasibility of such attacks on language models, showing that private information could be extracted through prompting with different strategies [57, 34, 58, 25, 59, 56].
The effectiveness of inference attacks can serve as a proxy measure for memorization. Models that are more susceptible to these attacks are generally considered to have higher levels of memorization.
In conclusion, while measuring memorization in language models remains a challenging task, techniques like the Exposure Metric and Inference Attacks provide valuable insights into the memorization behavior of these models. These methods allow researchers to approximate the extent of memorization and assess potential privacy risks associated with large language models.
III-C Counterfactual memorization
Previous work analyzed the memorization of large language models on sensitive information (e.g. phone numbers) in the training data [13] or synthetically injected ‘canaries’ [30, 60]. However, not all the memorized texts are equally interesting. Zhang et al. [61] dig into a new kind of memorization which is counterfactual memorization. The idea is to see how the presence or absence of a sample of the dataset, affects the performance of the model on the same sample. This definition of memorization has similarities with the definition of differential privacy. Differential privacy as mitigation strategy will be discussed in Section VI-C. In their experiments, the authors create different subsets of a bigger dataset and then they fine-tune the LM on each of these. Then they consider an item $x$ from the datasets and based on the presence or absence of $x$ in the subsets, they divide the subsets into two groups: IN and OUT. Then they test and report the performance on the IN and OUT group of models by averaging. Their experiments on 400 trained models show the counterfactual memorized data, are generally unconventional text such as all-caps, structured formats (i.e. tables or bullet lists), and multilingual texts. In conclusion, counterfactual memorization provides a novel perspective on understanding what language models retain from their training data.
IV Influences and dynamics of memorization
Understanding the factors that influence memorization in large language models is crucial for developing more efficient, secure, and privacy-preserving systems. This section explores various aspects that affect memorization, from model architecture to training processes.
IV-A Model capacity
The first significant factor influencing memorization is model size. Carlini et al. [23] and Tirumala et al. [31] have both demonstrated that larger models are more prone to memorization and do so more rapidly. This trend persists across different architectures and datasets. Carlini et al. [23] find that the relationship between model size and memorization grows consistently on a log-linear scale, with larger models memorizing a greater portion of the data.
Tirumala et al. [31] further highlight that larger models not only memorize more but do so faster in the training process. Interestingly, while memorization increases with model size, it does not necessarily correlate with improved performance. This was shown by Carlini et al. [23] by comparison of models with similar capacities but differing performance levels because of their architectures. Overall, the findings suggest that the ability of LLMs to memorize is strongly linked to their size, potentially due to the high capacity of these models to store detailed information from training data.
IV-B Training data characteristics
The nature of the training data heavily influences memorization. Lee et al. [62] develop tools to deduplicate training data and show that models trained on deduplicated data would produce memorized text ten times less frequently. Kandpal et al. [29] show that a sequence appearing 10 times in the training data is, on average, generated approximately 1000 times more frequently than a sequence that appears only once. A study done by Tirumala et al. [31] on memorization of different parts of speech, reveals that nouns and numbers are memorized significantly faster than other parts of speech, likely because they serve as unique identifiers for specific samples.
IV-C Input and prompting strategies
Carlini et al. [13], McCoy et al. [63], Kandpal et al. [29] show that longer prompts increase the likelihood of triggering memorized sequences, making it easier for language models to regurgitate training data. Moreover, methods like prefix tuning [64] and prompt engineering have been employed to maximize memorization. Ozdayi et al. [65] introduce a novel approach using prompt tuning to control memorized content extraction rates in LLMs. Weller et al. [66] propose ‘according-to’ prompting, a technique that directs LLMs to ground responses in previously observed text.
IV-D Tokenization
Kharitonov et al. [67] explore the impact of the tokenizer on memorization. They experiment with the size of the subword vocabulary learned through Byte-Pair Encoding (BPE) and demonstrate that increasing the subword vocabulary significantly affects the model’s ability and inclination to memorize training data. Furthermore, models with larger vocabularies are more likely to reproduce training data when given specific prompts. The authors suggest that this effect stems from the reduction in sequence lengths as BPE vocabulary size increases.
IV-E Sampling methods
In their experiments, Carlini et al. [13] initially opt for greedy sampling to maximize the regeneration of training data. One limitation is that this sampling scheme generates low diversity outputs; thus, they also experiment with the decaying temperature and Top-n sampling methods, the latter of which shows to be more successful. Yu et al. [68] experiment with different sampling schemes including Decaying temperature, Top-n, Nucleus-$\eta$ and Typical-$\phi$ samplings [69, 70, 71] and use an auto-tuning method on these to find the optimal sampling method that yields to the maximization of training data reproduction.
IV-F Fine-tuning and transfer learning
Mireshghallah et al. [72] evaluate how different fine-tuning methods — full model, model head, and adapter fine-tuning — vary in terms of memorization and vulnerability to privacy attacks. Their research, using membership inference and extraction attacks, finds that head fine-tuning is most susceptible to attacks, whereas adapter fine-tuning is less prone. Zeng et al. [73] conduct a comprehensive analysis of fine-tuning T5 models [74] across various tasks, including summarization, dialogue, question answering, and machine translation, finding that fine-tuned memorization varies significantly depending on the task. Additionally, they identify a strong link between attention score distributions and memorization, and propose that multi-task fine-tuning can mitigate memorization risks more effectively than single-task fine-tuning.
IV-G Training process dynamics
Zhang et al. [61], Kandpal et al. [29] show that memorization grows consistently with the number of training epochs, which makes sense, as more epochs push the model to potential overfitting. Jagielski et al. [75] demonstrate that examples seen during the earlier stages of training are less prone to memorization and rather they are forgotten over time. These findings indicate that memorization increases with more training, while early-seen examples being more likely to be forgotten.
IV-H Forgetting mechanisms
In machine learning, forgetting mechanisms are the processes through which models lose or discard previously learned information [76]. These mechanisms can occur unintentionally as part of the natural training dynamics or be purposefully induced to meet specific objectives, such as improving model generalization or addressing privacy concerns [77, 78, 79]. Blanco-Justicia et al. [80] provide a recent, detailed overview of forgetting in LLMs.
Catastrophic forgetting
Kirkpatrick et al. [81] initially introduced catastrophic forgetting in the context of neural networks and continual learning. They propose a method to protect important model weights to retain knowledge. This approach has been effective in maintaining performance on older tasks, even after long periods of non-use. Tirumala et al. [31] observe the forgetting mechanisms of a special batch through the learning process and show that it follows an exponentially degradation, reaching to a constant value baseline. They show that the mentioned baseline scales with the model size. Jagielski et al. [75] address the dual phenomena of memorization and forgetting in LLMs through stronger privacy attacks and several strategies for measuring the worst-case forgetting of the training examples. The study introduces a method to measure to what extent models forget specific training data, highlighting that standard image, speech, and language models do indeed forget examples over time, though non-convex models might retain data indefinitely in the worst case. The findings suggest that examples from early training phases, such as those used in pre-training large models, might enjoy privacy benefits but could disadvantage examples encountered later in training.
Intentional forgetting techniques
As memorization could lead to privacy risks and copyright issues, intentional forgetting could be necessary in some situations. Methods like knowledge unlearning aim to selectively remove specific information from trained models without retraining it from scratch. Bourtoule et al. [82] introduce the SISA framework for efficient machine unlearning, which divides the training data into shards (shards partition the data into disjoint segments) and trains sub-models that can be easily retrained if data needs to be removed. For LLMs specifically, Chen and Yang [83] introduce lightweight unlearning layers into transformers, allowing for selective data removal without full model retraining. Pawelczyk et al. [84] introduce “In-Context Unlearning”, which involves providing specific training instances with flipped labels and additional correctly labeled instances as inputs during inference, effectively removing the targeted information without updating model parameters. Additionally, knowledge unlearning techniques have been categorized into parameter optimization, parameter merging, and in-context learning, each offering unique advantages in efficiently removing harmful or undesirable knowledge from LLMs [85]. These methods not only enhance privacy and security but also ensure that the overall performance of the models remains intact, making them scalable and practical for real-world applications [86].
Factor from Section IV Key findings References
Model capacity Larger models memorize more Carlini et al. [23], Tirumala et al. [31]
Training data Duplicated data amplifies memorization Lee et al. [62], Kandpal et al. [29, 29], Tirumala et al. [31]
Input and prompting Longer prompts and prompt tuning can facilitate recall of the memorized suffix. Carlini et al. [13], McCoy et al. [63], Kandpal et al. [29], Li and Liang [64], Ozdayi et al. [65], Weller et al. [66]
Tokenization Bigger tokenizer vocabulary leads to more memorization Kharitonov et al. [67]
Sampling methods While greedy sampling can pinpoint extremely memorized samples, top-n sampling is the most effective method to retrieve more memorized items. Carlini et al. [13], Yu et al. [68]
Fine-tuning The amount of memorization after fine-tuning significantly varies depending on the task. Mireshghallah et al. [87], Zeng et al. [73]
Training process Earlier phases of training are less prone to memorization Zhang et al. [61], Kandpal et al. [29], Jagielski et al. [75]
Forgetting mechanisms Forgetting follows an exponentially decaying curve Tirumala et al. [31], Jagielski et al. [75]
TABLE II: Influences and dynamics of memorization and their key findings (discussed in Section IV)
In conclusion, memorization in LLMs is influenced by a complex interplay of factors ranging from model architecture to training dynamics. Understanding these influences is crucial for developing more robust, privacy-preserving, and ethically sound language models. An overview of the findings for each of the factors discussed in this subsections is provided in Table II.
V Memorization in specific model architectures
While the target of most of the papers discussed in this survey is auto-regressive large language models, in this section we will overview the memorization in other types of language models, as it would help us gain a deeper understanding of this phenomena.
V-A Masked language models (MLM)
Early studies investigating privacy leakage through the memorization of masked language models (MLMs) yielded limited success. For instance, Vakili and Dalianis [88] attempt to extract patients’ medical conditions from BERT models trained on clinical data, concluding that the risk of privacy leakage in BERT-based models is minimal. Similarly, Lehman et al. [89] conduct attacks on BERT models trained on the MIMIC-III corpus, reaching the same conclusion that stronger attacks may be necessary to trigger memorization in MLMs. Jagannatha et al. [90] employ membership inference attacks on clinical MLMs, demonstrating that these models are far less vulnerable to such attacks compared to auto-regressive (generative) models. Carlini et al. [23] extend their previous work on auto-regressive language models by conducting experiments on the T5 [74] model family, trained on the C4 corpus. They demonstrated that scaling laws still apply, but the degree of memorization in MLMs remains significantly lower compared to similarly sized auto-regressive models.
Mireshghallah et al. [87] introduce “likelihood ratio membership inference attacks” and show that privacy risks in MLMs, while smaller, can still be significant if the attacker uses more advanced methods.
V-B Aligned language models
In studies exploring memorization in aligned language models, experiments typically use either targeted prompting or random prompting approaches. However, these methods face two key challenges that limit their effectiveness.
First, the role tag which is prepended to both input queries and model responses disrupts traditional attack methods, such as the prefix prompting methods. Second, alignment training further complicates extraction attempts. Even if attackers could bypass the role tag, the model’s alignment training, designed to ensure ethical responses, reduces the likelihood of disclosing sensitive or inappropriate information.
To address these challenges, Nasr et al. [34] introduce the “Divergence Attack.” This method asks the model to endlessly repeat a specific word, causing it to break from its alignment training and revert to its original language modeling objectives. In some cases, this led to the verbatim reproduction of training data. Notably, this only occurs when single-token words were repeated, as multi-token prompts led to the model continuing the repetition without reproducing training data.
V-C Retrieval augmented generation (RAG)
Retrieval-augmented language models generate text distributions referencing both the parameters of the underlying language model and the information retrieved from a document index. Huang et al. [91] analyze privacy issues in retrieval-based language models. They use the kNN-LM [92] as their target model through API access. They investigate a scenario in which a model creator has a private, domain-specific data-store that improves model performance on domain-specific tasks, but may also contain sensitive information that should not be revealed. They analyze the access to the text completion and perplexity calculation APIs of the model. Their findings show that kNN-LMs are more prone to disclosing private data from their private data-store compared to parametric models. Zeng et al. [93] conduct targeted and prefix attacks on Retrieval-Augmented Generation (RAG) frameworks to examine how their memorization behaviors differ from those of standard LLMs. The results showed that incorporating retrieval mechanisms significantly lowers the model’s memorization capacity, reducing the success rate of training data extraction attacks in both targeted and prefix attack scenarios. This suggests that the use of retrieval-based data in RAG models makes them more resilient to such attacks compared to standalone LLMs, as the external retrieval component limits the model’s reliance on internal memorization.
VI Mitigating memorization: strategies and techniques
As discussed in earlier sections, memorization is influenced by a range of factors and it is sometimes necessary for the learning process [94]. However, in scenarios where memorization could lead to privacy concerns or security vulnerabilities, some methods could be employed to mitigate its impact. In these cases, specific strategies are utilized to limit the retention of sensitive information, ensuring that potential risks related to data exposure or misuse are minimized.
VI-A Data de-duplication
Lee et al. [62] run exact matching and approximate matching (MiniHash) de-duplication algorithms on the C4, RealNews, LM1B, and Wiki40B datasets and show that these datasets contain up to 13.6% near duplicates and up to 19.4% exact duplicates; showing that de-duplication can reduce training cost in the first place. Measuring the amount of duplicates shared between train and validation splits of these datasets, they show that there is a possibility of validation data leakage which causes the over-estimation of the model accuracy or may lead to biases when considering hyperparameters tuning.
To investigate the impact of data de-duplication on language model’s memorization, they train a 1.5B parameter GPT-2 model from scratch on three different settings: C4-Original, C4-NearDup, and C4-ExactSubstr, each for two epochs. Then the evaluate the memorization in no prompt and prompted settings and count the 50-token substrings generations that are presented in the training data as memorization. The no-prompt experiment generations show 10×10$\times$10 × less memorization in de-duplicated trained models. On the other hand, in the prompted experiment, when the prompt comes from the duplicate examples, the model trained on C4-Original generates the true exact continuation over 40% of the time. The other two models also generate the ground truth more often when the prompt is sampled from the duplicate examples, suggesting that more harsh de-duplication algorithms are needed to prevent memorization.
VI-B MemFree decoding
Ippolito et al. [32] try to prevent memorization using the novel MemFree Decoding method. Their approach proposes a new sampling strategy to prevent the emission of memorized samples. In the sampling phase, they keep track of the current n-gram (where ‘n‘ is a fixed parameter and should be chosen carefully; small n would prevent many common phrases to appear and big n may not prevent memorization) that is being generated and after generation of each token, they check the existence of the last n-gram in the training set. If this n-gram is a member of the training set, they ignore the newly generated token and sample another token, so as to obtain a non-memorized string. They inspect the GPT-Neo [95] family of models from smallest size to the largest one, prompting with 50 token sequences from the training data and anticipating for a 50 token suffix match. They show that their method significantly reduces the similarity of the generations with respect to the training data. They also show that although the approximate memorization increases with the model size log-linearly, MemFree defence stays effective in keeping the similarity around 0.6 BLEU score even in the largest model size.
While this verbatim memorization filtering reduces the average similarity score between the training data and the generations, qualitative analysis of the generations shows that the models opt to cheat in some sense. Ippolito et al. [32] inspect the production model GitHub Copilot [96] and show that when it is prompted with the training data, it successfully avoids verbatim generation, however, a simple style transfer in the prompts easily bypasses the Copilots memorization defence. They repeat this experiment on the English language models such as PaLM [36] and show how this defense could be bypassed, concluding this defence is an incomplete defence.
VI-C Differential privacy
Differential privacy (DP) is a technique that protects individual data points by adding noise, minimizing the impact of any single data point on the model’s output. DP-SGD (Differentially Private Stochastic Gradient Descent) is an adaptation of the standard SGD algorithm, designed to fine-tune language models while maintaining privacy [97]. Carlini et al. [30] demonstrate that by adjusting the privacy budget parameter $\epsilon$ in DP-SGD training, the exposure of memorized data can be reduced to a level that makes it indistinguishable from any other data. However, this comes at the cost of reduced model utility and a slower training process.
To address these utility issues, some studies propose selective differential privacy approaches [98, 99, 100, 101]. For instance, Kerrigan et al. [98] propose training a non-private base model on a public dataset and then fine-tuning it on a private dataset using DP-SGD. This approach aims to balance privacy with model performance.
Li et al. [99] show that with carefully chosen hyperparameters and downstream task objectives, fine-tuning pretrained language models with DP-SGD can yield strong performance on a variety of NLP tasks at privacy levels. Remarkably, some of their fine-tuned models even outperform non-private baselines and models trained under heuristic privacy approaches.
While DP is effective for reducing memorization and preserving privacy, it is crucial to select hyperparameters wisely. Otherwise, the model may not withstand stronger privacy attacks, potentially compromising its effectiveness as shown in Lukas et al. [59].
VI-D Unlearning methods
As discussed in the intentional forgetting section [IV-H], unlearning methods offer effective strategies to mitigate memorization by selectively removing specific information from trained models without the need for full retraining. However, a limitation of these unlearning methods is that they require prior knowledge of what the model has memorized in order to selectively remove it, which can be challenging to determine and may limit the effectiveness of these techniques.
VII Future directions and open challenges
Based on the discussion of the existing literature to date on memorization in LLMs, we make suggestions for research topics to be addressed in the near future.
VII-A Balancing performance with privacy in LLMs
As discussed earlier, memorization in LLMs can also have beneficial uses, such as improving factual recall and enhancing performance in specific tasks. The precondition is that privacy leakage and copyright issues have been taken care of. Methods like Differential Privacy (DP), that are designed to mitigate memorization risks and prevent sensitive data exposure, often come at the cost of reduced model performance, as they limit the model’s ability to retain useful information. To fully harness the benefits of memorization while safeguarding against privacy breaches, future research needs to focus on strategies that balance these competing goals. This includes developing techniques that protect sensitive data and intellectual property without significantly degrading the model’s accuracy and utility. Such efforts will be key to ensuring both ethical compliance and high-performance outcomes in LLMs.
VII-B Reducing factual memorization in favor of conceptual memorization
The interplay between factual and conceptual memorization is relevant in LLM development, because there is a trade-off between correctness of information, and undesirable recall of training data: While factual memorization can provide precise information recall, it also presents higher risks in terms of privacy and data protection. Conceptual memorization, on the other hand, contributes to the model’s ability to generalize and apply knowledge flexibly, but may be more challenging to audit and control. For example, when using LLMs for question answering, literal reproduction of facts is desired; this includes names and numbers (e.g. “Who was the first person to fly across the ocean and when did this take place?”). It is straightforward to evaluate LLMs for generating the correct facts. If we prefer conceptual memorization in LLMs, we should accept a higher rate of hallucinations, because the model more freely generates text based on its parametric knowledge.
Future research in this area should therefore focus on developing techniques to balance these two types of memorization, enhancing the benefits of each while mitigating their respective risks. This could involve methods to selectively encourage conceptual memorization while limiting unnecessary factual memorization, particularly of sensitive information, depending on the type of task at hand.
VII-C The boundary of memorization and understanding
It is relatively straightforward to design experiments that demonstrate a model’s ability to generate fluent and human-like text, such as generating novel content or adapting to new contexts. However, verifying that these capabilities are not simply the result of memorization rather than abstraction over the information processed during training is much more challenging, as both can produce similar outcomes. Distinguishing between these requires carefully designed experiments, and further research is needed to clarify when a model is abstracting over learned concepts versus merely memorizing and reproducing them.
VII-D Memorization in specific contexts
Several application domains are currently understudied with respect to the effect of memorization. We identified four contexts specifically where more research is needed.
Conversational Agents
LLMs that have been fine-tuned for conversations can be used as conversational agents (chatbots), e.g., to assist customers of online services [102, 103]. Nasr et al. [34] introduce their so-called “divergence attack” to extract memorized samples from conversation-aligned LLMs; however, as they also discuss, these attacks are not powerful enough to stimulate training data reproduction. This makes us conclude that attacking conversation-aligned language models requires more advanced methods. Since most of the production language models are only available in a conversational settings, addressing memorization in conversational agents and conducting novel attacking methods to alleviate training data extraction in these models would be a prominent research direction.
Retrieval-Augmented Generation (RAG)
In RAG frameworks, LLMs are helped with a retrieval component that first selects the most relevant documents from a collection and feeds them to the prompt [104, 105]. The LLM then generates an answer based on the provided sources. This has advantages such as the reduction of hallucination [106], the transparency of sources, and the potential of generating answers related to proprietary or novel sources that were not in the LLM training data. Although very popular, RAG is not yet thoroughly analyzed regarding memorization. This is important, because a reduction of hallucination when using RAG in LLMs could also have an increase in memorization as a side effect: less hallucination means that the LLM stays closer to the original content. Recently, Zeng et al. [93] analyzed privacy aspects in RAG. They show that RAG systems are vulnerable of leaking private information. On the other hand, they also found that RAG can mitigate the reproduction of the LLM training data. A broader study focusing on memorization in RAG-based LLMs that compares different proposed retrieval architectures is an important future research direction.
Multilingual Large Language Models (xLM)
xLMs are trained to interpret and generate text in multiple languages. Trained on extensive datasets that include various languages, these models develop language-agnostic representations [107]. Despite their potential, xLMs often face challenges related to data scarcity for low-resource languages, leading to performance disparities [108]. This necessitates ongoing research to enhance their efficacy and fairness across all languages. For future work, we propose to investigate whether a low-resource language setting is more prone to memorization by comparing to English scenarios. Addressing this research direction is essential not only for understanding the privacy implications associated with LLMs and data leakage in low-resource settings but also for ensuring AI safety in societies using LLMs with languages that have fewer resources.
Diffusion Language Models (DLM)
Since DLMs show remarkable performance in the vision domain, researchers have started to adopt the diffusion models (DM) idea to the text domain and utilize their generative capabilities [109]. Carlini et al. [110] have conducted data extraction analysis on the image diffusion models, showing that diffusion models are much less private than prior generative models such as GANs, and that mitigating these vulnerabilities may require new advances in privacy-preserving training. Gu et al. [111] also show that according to the training objective of the diffusion models, a memorization behavior is theoretically expected and then they quantify the impact of the influential factors on the memorization behaviors in DMs. However, research focusing on the memorization issues related to DLMs for text remains unexplored. Even though the idea of diffusion language models is the same as the vision domain, they are inherently different, because of the discrete nature of the text domain. Therefore an analysis of the general vision diffusion models may not be applicable to the diffusion language models, making an independent research on memorization against DLMs a prominent future direction.
VIII Conclusion
Figure 3: The distribution by publication year of the papers discussed in this survey
In this paper, we systematically organized, summarized, and discussed the existing scientific work related to unintended and undesirable memorization in LLMs. Undesirable memorization might lead to privacy risks and other ethical consequences. We found that there exists a large body of research on memorization in LLMs, given the young age of the technology: Transformer models were first developed in 2017 [112], and the first generative large language model with emergent abilities was released in 2022 [1, 5]. The majority of papers on the topic were therefore published in very recent years (see Figure 3). This indicates that the field is working fast to analyze memorization and developing methods to mitigate undesirable memorization.
Despite the fast-growing body of literature on the topic, we argue that there are important areas that require more attention in research in the coming years. We have pointed out four specific contexts in which memorization needs to be studied and, when needed and possible, mitigated: LLM-based conversational agents, retrieval-augmented generation, multilingual LLMs, and diffusion language models. These areas all are of high importance, not only from the academic perspective, but maybe even more from the application and industry perspective. In particular conversational agents and retrieval-augmented generation are actively being developed in the commercial context. When applied to proprietary databases or in interaction with customer information, these applications are particularly vulnerable to privacy and security risks.
Limitations
The main limitation of a SoK paper on a highly active research topic is the fast pace in which the field is moving. A survey paper written in 2024 risks to become outdated by 2026. Although we acknowledge this, we argue that the topic of memorization is too important to not create this systemization of knowledge – especially because of the added value that we created with our concrete suggestions for future directions.
Acknowledgment
This publication is part of the project LESSEN [^1] https://lessen-project.nl with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).